OpenStudy (anonymous):

A pair of linear equations is shown below: y = –x + 1 y = 2x + 4 Which of the following statements best explains the steps to solve the pair of equations graphically?

OpenStudy (anonymous):

On a graph, plot the line y = –x + 1, which has y-intercept = –1 and slope = 1, and y = 2x + 4, which has y-intercept = 2 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1, which has y-intercept = 1 and slope = 1, and y = 2x + 4, which has y-intercept = 1 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1, which has y-intercept = 1 and slope = –1, and y = 2x + 4, which has y-intercept = –2 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1, which has y-intercept = 1 and slope = –1, and y = 2x + 4, which has y-intercept = 4 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (anonymous):

On a graph, plot the line y = –x + 1, which has y-intercept = –1 and slope = 1, and y = 2x + 4, which has y-intercept = 2 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1, which has y-intercept = 1 and slope = 1, and y = 2x + 4, which has y-intercept = 1 and slope = 4, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1, which has y-intercept = 1 and slope = –1, and y = 2x + 4, which has y-intercept = –2 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution. On a graph, plot the line y = –x + 1, which has y-intercept = 1 and slope = –1, and y = 2x + 4, which has y-intercept = 4 and slope = 2, and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (anonymous):

sorry about the ��������������������������� it just means negative